A health clinic uses a solution of bleach to sterilize petri dishes in which cultures are grown. The sterilization tank contains 90 gal of a solution of 4% ordinary household bleach mixed with pure distilled water. New research indicates that the concentration of bleach should be 8% for complete sterilization. How much of the solution should be drained and replaced with bleach to increase the bleach content to the recommended level?
Answer by stanbon(75887)

A health clinic uses a solution of bleach to sterilize petri dishes in which cultures are grown. The sterilization tank contains 90 gal of a solution of 4% ordinary household bleach mixed with pure distilled water. New research indicates that the concentration of bleach should be 8% for complete sterilization. How much of the solution should be drained and replaced with bleach to increase the bleach content to the recommended level?
------
Equation:
bleach - bleach + bleach = bleach
----
0.04*90 - 0.04x + 1*x = 0.08*90
-----
Multiply thru by 100 to get:
4*90 - 4x + 100x = 8*90
96x = 4*90
x = (4*90)/96
x = 3.9167 gallons (amt. that should be removed and replaced)
===================
Cheers,
Stan H.